solve each system by substitution
5x - 2y = -5
y- 5x = 3

4x - 12y = 5
-x + 3y -1

Answer by rmromero(383)

5x - 2y = -5
y- 5x = 3
Here are the Steps using Substition method
1. In either equation, solve for one variable in terms of the other. Since y- 5x = 3 is simplier than the other eqution, we will solve x in terms of y.
y- 5x = 3
y = 5x + 3


2. Substitute for that variable in the other equation. Solve.
5x - 2y = -5 , y = 5x + 3
5x - 2(5x+3) = -5
5x -10x - 6 = -5
-5x = 1
x = -1/5


3. Substitute the result from step 2 in either equation. Solve for the other variable.
y- 5x = 3, x = -1/5
y - 5(-1/5) = 3
y + 1 = 3
y = 2


4. Check the solution in both original equations.
5x - 2y = -5, x = 2 and y = -1/5
5(-1/5) - 2(2) = -5
-1 - 4 = -5
-5 = -5 ----------->> True
y- 5x = 3
2 - 5(-1/5) = 3
2 + 1 = 3
3 = 3 ----------->> True


Therefore the solution of the systems of equation is x = -1/5 and y = 2



4x - 12y = 5
-x + 3y -1 <<<<<=== im not sure what do you mean by this. please encode it clearly.